OpenStudy (anonymous):

A lamppost casts a shadow that is 35 yards long. A 3-foot-tall mailbox casts a shadow that is 5 yards long. How tall is the lamppost?

OpenStudy (anonymous):

@goformit100 please help

OpenStudy (anonymous):

First, convert everything to feet.

OpenStudy (anonymous):

or convert everything to yards, for that matter.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

35/5=7, therefore the ratio between the height of the objects is 7

OpenStudy (anonymous):

7*3=24

OpenStudy (anonymous):

24*3=72 feet

OpenStudy (anonymous):

24 Yards or 72 Feet